Let be a real number. Let x~2.5 be an approximate value of with absolute error



at most 0.01. The function ( )





is evaluated at instead of . Estimate the absolute error


1
Expert's answer
2021-12-31T08:56:19-0500

for example "f(x)=x^2"

then absolute error of f(x) at x = 2.5:

"|f(x+\Delta x)-f(x)|=f(2.51)-f(2.5)=2.51^2-2.5^2=0.0501"